(D^2-D'^2-3D+3D')z=e^x-2y


1
Expert's answer
2021-05-07T12:24:12-0400

There's some typing mistake in question, the correct equation is

"(D^2-3D)y=e^x-2y"

"\Rightarrow(D^2-3D+2)y=e^x"

Auxiliary Equation (A.E.) is "D^2-3D+2=0"

"\Rightarrow D^2-2D-D+2=0"

"\Rightarrow D(D-2)-(D-2)=0"

"\Rightarrow (D-1)(D-2)=0"

"\therefore \space D=1,2"


Complimentary Function (C.F) is

"y=c_1e^x+c_2e^{2x}"


For Particular Integral (P.I)

P.I "=\dfrac{1}{D^2-3D+2}e^x"

"=\dfrac{1}{(D-1)(D-2)}e^x"

"=-1\bigg[\dfrac{1}{(D-1)}e^x\bigg]"

"=-1.x.\dfrac{1}{1}e^x"

"=-xe^x"


Complete Solution (C.S) = C.F + P.I

"y=c_1e^x+c_2e^{2x}-xe^x"